OpenStudy (timaashorty):

Are the given vectors normal? a=(5,-2) and b=(6,15)

OpenStudy (timaashorty):

@tHe_FiZiCx99

OpenStudy (anonymous):

i think you only need to check if the dot product is zero, because "normal" in this case means "perpendicular"

OpenStudy (anonymous):

\[<5,-2>\cdot<6,15>=5\times 6+(-2)\times 15\] if you get zero, then "normal"

OpenStudy (anonymous):

which of course you do in this case since it is \(30-30=0\)

OpenStudy (timaashorty):

Omg had no idea you responded. Thank you! So what you're basically saying is that it is normal since it equals zero?

OpenStudy (anonymous):

since the dot product is zero, then they are normal

OpenStudy (timaashorty):

now I see. Thank you so much!

OpenStudy (anonymous):

yw